40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 9 23:18:09 2012, yuta, Update, LSC, PRMI got more stable a bit LSCMICHOLTF.pngLSCPRCLOLTF.pngALL_1025928219_PRMIlocked3.pngPRMIbeammotion20120709.png
    Reply  Tue Jul 10 01:52:47 2012, Koji, Update, LSC, PRMI got more stable a bit 
       Reply  Tue Jul 10 03:16:17 2012, yuta, Update, LSC, PRMI got more stable a bit LSCMICHOLTF.pngLSCPRCLOLTF.pngPRMI20120709_2.png
Message ID: 6950     Entry time: Tue Jul 10 03:16:17 2012     In reply to: 6949
Author: yuta 
Type: Update 
Category: LSC 
Subject: PRMI got more stable a bit 

I modified filiters for LSC_MICH and LSC_PRCL again to cope with power recycling gain fluctuation.
After some more alignment, power recycling gain increased (but still ~3.7). It fluctuates more than a factor of 2, and I began to see glitches again. So I needed more gain margin, as Koji pointed out.

I played around with filters, but I couldn't remove all the glitches. Gain margin now look OK in principle.
It looks like PRM motion is related. Since PRM doesn't have oplev now, I will see PRM oplev tomorrow.

New openloop transfer function:
 LSC_MICH
   UGF ~100 Hz, phase margin ~ 50 deg
   no phase flip in less than factor of ~5 gain change
   550 usec delay
 LSC_PRCL
   UGF ~100 Hz, phase margin ~ 70 deg (phase bump at UGF)
   no phase flip in less than factor of ~5 gain change
   550 usec delay
LSCMICHOLTF.pngLSCPRCLOLTF.png

Power recylcing gain:
  It is now ~3.7. It fluctuates pretty much. See time series data below when I locked PRMI. MICH and PRCL locks at the same time.

G = (1600-244)/(266-244)*0.06 = 3.7

PRMI20120709_2.png
 

ELOG V3.1.3-